OpenStudy (anonymous):

Multiply these binomials. (8x + 4)(x + 8) =

OpenStudy (anonymous):

8x^2+64x+4x+32 8x^2+68x+32

rvc (rvc):

a+b(c+d)=a(c+d)+b(c+d) use this @tempers_the_name

OpenStudy (whpalmer4):

\[(a+b)(c+d) = a(c+d) + b(c+d) = a*c + a*d + b*c + b*d\]